Find the focus and directrix of the parabola y^2=6x
Answer by morganjd(5)

y^2=6x
So The equation y^2=6x has the form y^2 = 4px, with 4px = 6, so p = 2/3.. The parabola (p) is horizontal, with FOCUS: ( 2/3, 0), directrix x = -2/3.
I hope this helps.
jdm